OpenStudy (cutiecomittee123):

Help me Solve this system of equations to get the sum of all of the x coordinates from the solutions. x^2+4y^2=100 4y-x^2=-20

OpenStudy (anonymous):

you posted this question earlier

OpenStudy (cutiecomittee123):

Im pretty sure it was involving different equations

OpenStudy (anonymous):

The one from before were different equations. But the same idea holds. You want to make some sort of substitution that will let you reduce one of the equations to only x or only y. So you're looking for something in common in both equations that will make for an easy substitution. I see that both equations have an x^2 term. So if I were to solve the first equation for x^2, I would get x^2 = 100 - 4y^2 With me so far?

OpenStudy (cutiecomittee123):

Yes I am with ya

OpenStudy (cutiecomittee123):

wait wouldnt you get 4y^2+4y=80 because you can cancel out the two x^2's since one is negative and if you add the equations it works right?

OpenStudy (anonymous):

Well, if we do the substitution for x^2 into the 2nd equation. we get 4y - (100-4y^2) = -20 4y - 100 + 4y^2 = -20 4y^2 + 4y = 80 SO yes, you're right, good job :) Would you know what to do from here?

OpenStudy (cutiecomittee123):

Now move 80 to the other side and you get 4y^2+4y-80=0 Then factor

OpenStudy (cutiecomittee123):

thats where I get confused because what adds to get 4 but multiplies to get -80???

OpenStudy (anonymous):

Well, conveniently all of those terms are divisible by 4, so let's just divide out the 4 from all terms. Doing this would reduce the quadratic into y^2 + y - 20 = 0

OpenStudy (cutiecomittee123):

thats a lot cleaner and easier to deal with

OpenStudy (cutiecomittee123):

now what

OpenStudy (anonymous):

Well, how does that factor?

OpenStudy (cutiecomittee123):

IDK???

OpenStudy (anonymous):

Well, -20 would need to factor into 2 numbers that would also add up to a positive 1. This means: One of the numbers has to be positive and the other has to be negative (otherwise they would never multiply to a negative 20) The largest of those 2 numbers must be positive. So can you come up with the two numbers? Two factors that multiply to -20 yet add to positive 1

OpenStudy (cutiecomittee123):

no, I bet its really easy but I just cant wrap my head around it.

OpenStudy (anonymous):

It is easy. And youll think its easy also once you get used to the process. Okay, so let's look at how 20 could possibly break apart: